### v2

# Geometry is Wavy: Curvature Wave Equations for Generic Affine Connections

Abstract
Geometry is wavy: even at the purely geometric level (no particular theory chosen), curvature satisfies a covariant quasilinear wave equation. In Riemannian geometry equipped with the Levi-Civita connection, the Riemann curvature tensor obeys a wave equation of the schematic form
$\Box Riem=\mathcal{Q}(Riem,Riem),$
where $\mathcal{Q}(Riem,Riem)$ denotes the terms quadratic in the curvature arising from the Bianchi identities. In this work, we generalize this curvature wave equation to spacetimes endowed with a generic affine connection possessing torsion and nonmetricity. Working within the metric-affine framework, we derive the corresponding wave equation for the Riemann tensor and analyze its structure in several geometrically and physically distinguished settings, including Einstein spaces, teleparallel gravity, and Einstein-Cartan theory.
Dedicated to Metin Gürses on the occasion of his becoming an emeritus professor after 60 years of research and teaching.
I INTRODUCTION
In her beautiful and compact book [1], Choquet–Bruhat shows that the Riemann curvature tensor, in the standard geometric setting of a torsion-free and metric-compatible connection, satisfies a nonlinear second-order wave equation of the form
$\nabla^{\alpha}\nabla_{\alpha}R_{\beta\gamma\lambda\mu}+S_{\beta\gamma\lambda\mu}+\Big\{\nabla_{\gamma}\big(\nabla_{\lambda}R_{\mu\beta}-\nabla_{\mu}R_{\lambda\beta}\big)-(\beta\to\gamma)\Big\}=0,$ (1)
where the quadratic term in the curvature is given as
$S_{\beta\gamma\lambda\mu}$ $=\Big\{R_{\gamma}{}^{\rho}\,R_{\rho\beta\lambda\mu}+R^{\alpha}{}_{\gamma}{}^{\rho}{}_{\beta}\,R_{\alpha\rho\lambda\mu}+\Big[\,R^{\alpha}{}_{\gamma}{}^{\rho}{}_{\lambda}\,R_{\alpha\beta\rho\mu}-(\lambda\to\mu)\Big]\Big\}-(\beta\to\gamma).$ (2)
This remarkable equation is purely geometric: it holds in dimensions greater than three without invoking any field equations, and follows solely from the differential Bianchi identities. In three dimensions, however, the differential Bianchi identity does not carry independent content beyond the contracted one; consequently, the curvature-wave equation does not arise. The metric signature also plays no essential role at this stage: in Riemannian signature, the same equation becomes a nonlinear elliptic equation rather than a hyperbolic one.
An important feature of (1) is that its trace does not yield a nontrivial equation; in particular, the Ricci tensor does not satisfy a closed wave equation on its own. Nevertheless, once the geometry is assumed to obey the Einstein field equations,
$G_{\alpha\beta}+\Lambda g_{\alpha\beta}=\kappa\,T_{\alpha\beta},$ (3)
with $G_{\alpha\beta}:=R_{\alpha\beta}-\tfrac{1}{2}g_{\alpha\beta}R$ and $\kappa=8\pi G$ (in units $c=1$), the Ricci tensor becomes algebraically determined by the energy-momentum tensor:
$R_{\alpha\beta}=\kappa\left(T_{\alpha\beta}-\frac{1}{2}g_{\alpha\beta}T\right)+\Lambda g_{\alpha\beta}.$ (4)
Introducing the trace-reversed energy-momentum tensor
$\widetilde{T}_{\alpha\beta}:=T_{\alpha\beta}-\frac{1}{2}g_{\alpha\beta}T,$ (5)
this relation takes the compact form
$R_{\alpha\beta}=\kappa\,\widetilde{T}_{\alpha\beta}+\Lambda g_{\alpha\beta}.$ (6)
Substituting this into (1), the curvature-wave equation reduces to
$\nabla^{\alpha}\nabla_{\alpha}R_{\beta\gamma\lambda\mu}+S_{\beta\gamma\lambda\mu}+\kappa\Big\{\nabla_{\gamma}\big(\nabla_{\lambda}\widetilde{T}_{\mu\beta}-\nabla_{\mu}\widetilde{T}_{\lambda\beta}\big)-(\beta\to\gamma)\Big\}=0.$ (7)
In a vacuum, the source terms in the curly brackets vanish identically; the first two terms survive and constitute a quasilinear wave equation satisfied by the Riemann tensor.
While it is not entirely clear who first derived this curvature-wave equation in the history of differential geometry, an early reference is [2], where it appears under the name of the Penrose wave equation and is employed to provide an elegant derivation of the Teukolsky equation governing perturbations of the Kerr geometry. More recently, a linearized version of this equation was used in [3] to analyze the gravitational wave memory effect. Despite these appearances, we believe that the wave equation obeyed by the Riemann curvature tensor deserves much wider attention, as it directly encodes the intrinsically wavelike character of geometry in Lorentzian signature and its elliptic character in Euclidean signature.
In this work, we extend this discussion beyond the Riemannian setting by allowing for a general affine connection that is neither symmetric nor metric-compatible. Although the resulting computations are considerably more involved, the geometric and physical insights gained from the general result justify the effort. Before presenting these results, we briefly review the relevant geometric frameworks. Our conventions are those of [4].
Einstein’s general theory of relativity [5] describes gravitation as a manifestation of spacetime curvature, encoded in the Riemann tensor constructed from the Levi-Civita connection of a Lorentzian metric. The differential geometry used in Einstein’s theory goes back to [6] and [7]. This formulation rests on two assumptions: the absence of torsion and the metric compatibility of the affine connection. While natural and economical, these assumptions are not imposed by first principles. From a differential-geometric viewpoint, the most general affine connection on a manifold may possess torsion and nonmetricity in addition to curvature.
Relaxing these assumptions leads to the framework of metric-affine geometry, where the metric and affine connection are treated as independent structures. In this setting, gravitational interactions may be encoded not only in curvature but also in torsion and nonmetricity. Torsion measures the antisymmetric part of the connection and governs the failure of infinitesimal parallelograms to close, while nonmetricity quantifies the variation of the metric under parallel transport. Together with curvature, these tensors provide a complete local characterization of general affine geometry.
Several important gravitational theories arise as consistent restrictions of the metric-affine framework. General relativity is recovered when both torsion and nonmetricity vanish. Einstein–Cartan theory [8] allows for nonvanishing torsion while preserving metric compatibility, thereby coupling spacetime geometry to the intrinsic spin of matter [9]. Teleparallel theories [10, 11], by contrast, are characterized by vanishing curvature, with gravitation encoded entirely in torsion and/or nonmetricity. In particular, metric teleparallel gravity employs torsion alone [10], symmetric teleparallel gravity employs nonmetricity alone [12, 13, 14, 15], and more general teleparallel models allow for both.
These interrelations are most transparently revealed through the decomposition of a general affine connection into its Levi-Civita [7] part and tensorial corrections associated with torsion and nonmetricity. This geometric landscape shows that gravity may be equivalently described in terms of curvature, torsion, or nonmetricity, depending on the chosen representation. See Figure 1 for a classification of these theories. [^1] We owe the last line of the table to F.W. Hehl, see the nice exposition [16]
Adopting the metric-affine perspective, we derive the wave equation satisfied by the Riemann curvature tensor in the presence of both torsion and nonmetricity. We then specialize our general results to several geometrically and physically distinguished settings, including Einstein spaces, teleparallel gravity [10], and Einstein–Cartan theory [8]. This work highlights how additional geometric structures modify the propagation and dynamics of spacetime curvature, and it is dedicated to Metin Gürses on the occasion of his becoming an emeritus professor in recognition of his profound contributions to the geometry of gravitation.
Outline of the paper: In Section II we fix conventions and summarize the metric-affine geometric ingredients used throughout (torsion, nonmetricity, and the decomposition of a generic connection into Levi-Civita plus distortion). Section III derives the curvature, the Ricci tensor, and the scalar curvature in terms of these fields and discusses which familiar Riemannian symmetries survive in the generic case. In Section IV we establish the generalized Ricci identities and Bianchi identities needed for consistency checks and for manipulating covariant derivatives. Section V presents the field equations and the main results of the paper, together with the limits that recover standard general relativity. The appendices collect technical derivations and complementary material: Appendix A introduces the decomposition of the generic affine connection into the Levi Civita connection, torsion, and nonmetricity contributions. Appendix B gives the decomposition of the curvature tensors. Appendix C provides the commutator (Ricci) identities; Appendix D derives the generalized Bianchi identities; Appendix E discusses the teleparallel and symmetric-teleparallel specializations [10], while Appendix F summarizes the tetrad and spin connection [9] formulation and its relation to the metric-affine variables. The last appendix is devoted to a short historical survey.
Figure 1: Schematic classification of gravitational theories according to the geometric properties of spacetime. The table displays the presence or absence of curvature (Riemann tensor $R^{\rho}{}_{\mu\sigma\nu}$), torsion ($T^{\sigma}{}_{\mu\nu}$), and nonmetricity ($Q_{\mu\nu\sigma}$), together with the corresponding form of the affine connection $\Gamma^{\sigma}{}_{\mu\nu}$. Here $^{c}\Gamma^{\sigma}{}_{\mu\nu}$ denotes the Levi–Civita connection, $K^{\sigma}{}_{\mu\nu}$ the contorsion tensor associated with torsion, and $L^{\sigma}{}_{\mu\nu}$ the disformation tensor associated with nonmetricity. General relativity is recovered when both torsion and nonmetricity vanish. Einstein–Cartan theory allows for nonvanishing torsion while preserving metric compatibility. Teleparallel theories are characterized by the vanishing of the curvature, with gravitation encoded entirely in torsion and/or nonmetricity.
II CONVENTIONS
In a metric-affine spacetime, the metric $g_{\mu\nu}$ and the affine connection $\Gamma^{\sigma}{}_{\mu\nu}$ are treated as independent geometric objects. The departure from Riemannian geometry is encoded in two tensorial quantities: the nonmetricity
$Q_{\mu\nu\sigma}:=\nabla_{\mu}g_{\nu\sigma},$ (8)
which measures the failure of the connection to preserve the metric under parallel transport and the torsion
$T^{\sigma}{}_{\mu\nu}:=\Gamma^{\sigma}{}_{\mu\nu}-\Gamma^{\sigma}{}_{\nu\mu}$, (9)
which captures the antisymmetric part of the connection and governs the nonclosure of infinitesimal parallelograms.
[…]
$\Gamma^{\sigma}{}_{\mu\nu}={}^{c}\Gamma^{\sigma}{}_{\mu\nu}+K^{\sigma}{}_{\mu\nu}+L^{\sigma}{}_{\mu\nu}$. (10)
Here $^{c}\Gamma^{\sigma}{}_{\mu\nu}$ denotes the Levi-Civita (Christoffel) connection [7],
$^{c}\Gamma^{\sigma}{}_{\mu\nu}=\frac{1}{2}\,g^{\sigma\lambda}\big(\partial_{\mu}g_{\nu\lambda}+\partial_{\nu}g_{\mu\lambda}-\partial_{\lambda}g_{\mu\nu}\big),$ (11)
$K^{\sigma}{}_{\mu\nu}$ is the contorsion tensor constructed from the torsion,
[…]
and $L^{\sigma}{}_{\mu\nu}$ is the disformation tensor defined in terms of the nonmetricity. With our index conventions, a convenient and standard choice is [^3] This definition ensures that $L^{\sigma}{}_{\mu\nu}$ is symmetric in $\mu\nu$ whenever $Q_{\lambda\mu\nu}$ is symmetric in $\mu\nu$. Adopting a different sign convention for $Q_{\mu\nu\sigma}$ would correspondingly modify the signs in $L^{\sigma}{}_{\mu\nu}$.
[…]
Lowering the first index with the metric, these tensors obey useful symmetry properties
$Q_{\sigma\mu\nu}$ $=Q_{\sigma\nu\mu},$ $L_{\sigma\mu\nu}$ $=L_{\sigma\nu\mu},$ (14)
together with the antisymmetry of the contorsion in its first two indices,
$K_{\sigma\mu\nu}=-K_{\mu\sigma\nu}.$ (15)
Moreover, the contorsion tensor is related to the torsion through the identities
$K_{\mu\sigma\nu}-K_{\nu\sigma\mu}$ $=T_{\sigma\nu\mu},$ (16)
$K_{\mu\nu\sigma}-K_{\mu\sigma\nu}$ $=T_{\mu\nu\sigma}.$ (17)
Throughout this work, the covariant derivative $\nabla_{\mu}$ is defined using the full affine connection $\Gamma^{\sigma}{}_{\mu\nu}$. In particular,
$\nabla_{\mu}g_{\nu\sigma}=\partial_{\mu}g_{\nu\sigma}-\Gamma^{\lambda}{}_{\nu\mu}g_{\lambda\sigma}-\Gamma^{\lambda}{}_{\sigma\mu}g_{\nu\lambda},$ (18)
which reproduces the definition of the nonmetricity tensor (8).
III THE CURVATURE TENSORS AND THE BIANCHI IDENTITIES
The Riemann tensor [6] of the generic connection is defined by
$R^{\rho}{}_{\mu\sigma\nu}:=\partial_{\sigma}\Gamma^{\rho}{}_{\mu\nu}-\partial_{\nu}\Gamma^{\rho}{}_{\mu\sigma}+\Gamma^{\rho}{}_{\lambda\sigma}\Gamma^{\lambda}{}_{\mu\nu}-\Gamma^{\rho}{}_{\lambda\nu}\Gamma^{\lambda}{}_{\mu\sigma}$, (19)
and remains antisymmetric in its last two indices,
$R^{\rho}{}_{\mu\sigma\nu}=-R^{\rho}{}_{\mu\nu\sigma}$. (20)
In terms of the decomposition (10), one can express $R^{\rho}{}_{\mu\sigma\nu}$ as the Levi-Civita [7] curvature plus terms involving $K^{\sigma}{}_{\mu\nu}$ and $L^{\sigma}{}_{\mu\nu}$. (A detailed derivation is provided in Appendix B.) We denote the Levi-Civita curvature by $^{c}R^{\rho}{}_{\mu\sigma\nu}$. The Ricci tensor and scalar curvature are defined by the standard contractions
$R_{\mu\nu}:=R^{\rho}{}_{\mu\rho\nu},\qquad R:=g^{\mu\nu}R_{\mu\nu},$ (21)
and they can be written in terms of $^{c}R_{\mu\nu}$, $K^{\sigma}{}_{\mu\nu}$ and $L^{\sigma}{}_{\mu\nu}$ as recorded in Appendix B. A key technical input is the Ricci identity for a general connection acting on an arbitrary tensor $T^{\sigma\rho\cdots}{}_{\lambda\kappa\cdots}$: [^4] A proof is given in Appendix C.
$\big[\nabla_{\mu},\nabla_{\nu}\big]T^{\sigma\rho\cdots}{}_{\lambda\kappa\cdots}$ $=T^{\gamma}{}_{\mu\nu}\,\nabla_{\gamma}T^{\sigma\rho\cdots}{}_{\lambda\kappa\cdots}-\;R^{\sigma}{}_{\gamma\mu\nu}\,T^{\gamma\rho\cdots}{}_{\lambda\kappa\cdots}-\;R^{\rho}{}_{\gamma\mu\nu}\,T^{\sigma\gamma\cdots}{}_{\lambda\kappa\cdots}$-$\cdots$
$\quad+\;R^{\gamma}{}_{\lambda\mu\nu}\,T^{\sigma\rho\cdots}{}_{\gamma\kappa\cdots}+\;R^{\gamma}{}_{\kappa\mu\nu}\,T^{\sigma\rho\cdots}{}_{\lambda\gamma\cdots}$$\cdots$. (22)
The first and second Bianchi identities take the forms  [^5] Detailed computations are given in Appendix D.
$R^{\lambda}{}_{\nu\rho\sigma}+R^{\lambda}{}_{\sigma\nu\rho}+R^{\lambda}{}_{\rho\sigma\nu}$ $-\nabla_{\nu}T^{\lambda}{}_{\sigma\rho}-\nabla_{\sigma}T^{\lambda}{}_{\rho\nu}-\nabla_{\rho}T^{\lambda}{}_{\nu\sigma}$
$\quad+T^{\gamma}{}_{\sigma\rho}T^{\lambda}{}_{\gamma\nu}+T^{\gamma}{}_{\rho\nu}T^{\lambda}{}_{\gamma\sigma}+T^{\gamma}{}_{\nu\sigma}T^{\lambda}{}_{\gamma\rho}=0,$ (23)
$\nabla_{\nu}R^{\mu}{}_{\lambda\rho\sigma}+\nabla_{\rho}R^{\mu}{}_{\lambda\sigma\nu}+\nabla_{\sigma}R^{\mu}{}_{\lambda\nu\rho}$ $+R^{\mu}{}_{\lambda\nu\gamma}T^{\gamma}{}_{\rho\sigma}+R^{\mu}{}_{\lambda\rho\gamma}T^{\gamma}{}_{\sigma\nu}+R^{\mu}{}_{\lambda\sigma\gamma}T^{\gamma}{}_{\nu\rho}=0.$ (24)
A crucial qualitative difference from the Riemannian case is that, in general, one loses the symmetry under the exchange of index pairs,
$R_{\mu\nu\rho\sigma}\neq R_{\rho\sigma\mu\nu},$ (25)
because metric compatibility and torsion-freeness are no longer assumed.
IV A COMPACT FORM OF THE SECOND BIANCHI IDENTITY AND ITS TRACE
For later convenience, we introduce the torsion-dependent and nonmetricity-dependent combinations
$X^{\mu}{}_{\lambda\nu\rho\sigma}$ $:=R^{\mu}{}_{\lambda\nu\gamma}T^{\gamma}{}_{\rho\sigma}+R^{\mu}{}_{\lambda\rho\gamma}T^{\gamma}{}_{\sigma\nu}+R^{\mu}{}_{\lambda\sigma\gamma}T^{\gamma}{}_{\nu\rho}$, (26)
$Y_{\gamma\lambda\nu\rho\sigma}$ $:=-\,R^{\mu}{}_{\lambda\rho\sigma}\,Q_{\nu\mu\gamma}-\,R^{\mu}{}_{\lambda\sigma\nu}\,Q_{\rho\mu\gamma}-\,R^{\mu}{}_{\lambda\nu\rho}\,Q_{\sigma\mu\gamma}.$ (27)
Then, the second Bianchi identity 24 can be written in a compact form
$\nabla_{\nu}R_{\gamma\lambda\rho\sigma}+\nabla_{\rho}R_{\gamma\lambda\sigma\nu}+\nabla_{\sigma}R_{\gamma\lambda\nu\rho}+X_{\gamma\lambda\nu\rho\sigma}+Y_{\gamma\lambda\nu\rho\sigma}=0.$ (28)
As expected, $X_{\gamma\lambda\nu\rho\sigma}=0$ when torsion vanishes, and $Y_{\gamma\lambda\nu\rho\sigma}=0$ when the connection is metric compatible. Tracing (28) with $g^{\nu\gamma}$ yields a generalized contracted Bianchi identity. After carefully keeping track of the nonmetricity terms arising from raising indices, one arrives at
$\nabla^{\gamma}R_{\gamma\lambda\rho\sigma}+\nabla_{\sigma}R_{\lambda\rho}-\nabla_{\rho}R_{\lambda\sigma}+\tilde{Y}_{\lambda\rho\sigma}+X_{\gamma\lambda}{}^{\gamma}{}_{\rho\sigma}=0,$ (29)
where we have defined
$\tilde{Y}_{\lambda\rho\sigma}:=Y_{\gamma\lambda}{}^{\gamma}{}_{\rho\sigma}-R_{\gamma\lambda\sigma\nu}\,Q_{\rho}{}^{\nu\gamma}-R_{\gamma\lambda\nu\rho}\,Q_{\sigma}{}^{\nu\gamma}$. (30)
Similarly, it is convenient to rewrite the first Bianchi identity (23) as
$R^{\lambda}{}_{\nu\rho\sigma}+R^{\lambda}{}_{\sigma\nu\rho}+R^{\lambda}{}_{\rho\sigma\nu}+Z^{\lambda}{}_{\nu\rho\sigma}=0,$ (31)
where
$Z^{\lambda}{}_{\nu\rho\sigma}$ $:=-\nabla_{\nu}T^{\lambda}{}_{\sigma\rho}-\nabla_{\sigma}T^{\lambda}{}_{\rho\nu}-\nabla_{\rho}T^{\lambda}{}_{\nu\sigma}+T^{\gamma}{}_{\sigma\rho}T^{\lambda}{}_{\gamma\nu}+T^{\gamma}{}_{\rho\nu}T^{\lambda}{}_{\gamma\sigma}+T^{\gamma}{}_{\nu\sigma}T^{\lambda}{}_{\gamma\rho}$. (32)
V WAVE-TYPE EQUATION FOR THE RIEMANN TENSOR
We now derive a wave-type equation for the curvature in the presence of torsion and nonmetricity. Taking $\nabla^{\nu}$ of (28) and rearranging terms gives
$\nabla^{\nu}\nabla_{\nu}R_{\gamma\lambda\rho\sigma}+\nabla^{\nu}\nabla_{\rho}R_{\gamma\lambda\sigma\nu}+\nabla^{\nu}\nabla_{\sigma}R_{\gamma\lambda\nu\rho}+\nabla^{\nu}\!\left(X_{\gamma\lambda\nu\rho\sigma}+Y_{\gamma\lambda\nu\rho\sigma}\right)=0.$ (33)
Introducing $\Box:=\nabla^{\nu}\nabla_{\nu}$ and commuting derivatives in the second and third terms yield
$\Box R_{\gamma\lambda\rho\sigma}+\big[\nabla^{\nu},\nabla_{\rho}\big]R_{\gamma\lambda\sigma\nu}-\big[\nabla^{\nu},\nabla_{\sigma}\big]R_{\gamma\lambda\rho\nu}+\nabla_{\rho}\nabla^{\nu}R_{\gamma\lambda\sigma\nu}$
$\quad\quad-\nabla_{\sigma}\nabla^{\nu}R_{\gamma\lambda\rho\nu}+\nabla^{\nu}\!\left(X_{\gamma\lambda\nu\rho\sigma}+Y_{\gamma\lambda\nu\rho\sigma}\right)=0.$ (34)
At this point, one may attempt to simplify the last-index divergences $\nabla^{\nu}R_{\gamma\lambda\sigma\nu}$ and $\nabla^{\nu}R_{\gamma\lambda\rho\nu}$ using (28) and (29). In a generic metric-affine geometry, however, these terms do not collapse as neatly as in the Levi-Civita case; this is one of the technical distinctions that makes the metric-affine curvature wave equation structurally richer.
To proceed systematically, we evaluate the commutators in (34) using the Ricci identity (22). After a straightforward but lengthy computation, one arrives at the following wave-type equation (which is our main equation):
$\Box R_{\gamma\lambda\rho\sigma}+\nabla_{\rho}\nabla^{\nu}R_{\gamma\lambda\sigma\nu}-\nabla_{\sigma}\nabla^{\nu}R_{\gamma\lambda\rho\nu}+\nabla^{\nu}\!\left(X_{\gamma\lambda\nu\rho\sigma}+Y_{\gamma\lambda\nu\rho\sigma}\right)$ $+T^{\nu k}{}_{\rho}\,\nabla_{\nu}R_{\gamma\lambda\sigma k}-T^{\nu k}{}_{\sigma}\,\nabla_{\nu}R_{\gamma\lambda\rho k}$ $+R^{k}{}_{\gamma\rho}{}^{\nu}R_{k\lambda\sigma\nu}+R^{k}{}_{\lambda\rho}{}^{\nu}R_{\gamma k\sigma\nu}+R^{k}{}_{\sigma\rho}{}^{\nu}R_{\gamma\lambda k\nu}+R^{\kappa}\ _{\rho}R_{\gamma\lambda\sigma\kappa}$ $-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}-R^{\kappa}\ _{\sigma}R_{\gamma\lambda\rho\kappa}$ $=0,$ (35)
where $R^{k}{}_{\ \rho}:=R^{k}{}_{\nu\rho}{}^{\nu}=R_{\rho}{}^{k}$ denotes the Ricci tensor with one index raised. Equation (35) is the metric-affine generalization of the familiar curvature wave equation in Riemannian geometry: it reduces to the standard Levi–Civita result when $T^{\sigma}{}_{\mu\nu}=0$ and $Q_{\mu\nu\sigma}=0$, so that $X_{\gamma\lambda\nu\rho\sigma}=0=Y_{\gamma\lambda\nu\rho\sigma}$ and the additional torsion-driven transport terms vanish. Finally, for completeness, we explicitly record the dependence of $X$ and $Y$ on torsion and nonmetricity:
[…]
$Y_{\gamma\lambda\nu\rho\sigma}$ $=-\,R^{\mu}{}_{\lambda\rho\sigma}Q_{\nu\mu\gamma}-\,R^{\mu}{}_{\lambda\sigma\nu}Q_{\rho\mu\gamma}-\,R^{\mu}{}_{\lambda\nu\rho}Q_{\sigma\mu\gamma}.$ (37)
In the next section, we analyze several instructive special cases (Einstein spaces, teleparallel gravity [10], and Einstein-Cartan theory [8]), thereby illustrating how torsion and nonmetricity affect the propagation of curvature. [^6] Teleparallel gravity and its subclasses are discussed in Appendix E.
V.1 Wave equation in Einstein spaces
In this subsection, we specialize the general curvature wave equation (35) to Einstein spaces, i.e., spacetimes whose Ricci tensor is proportional to the metric:
$R_{\mu\nu}=\lambda\,g_{\mu\nu},$ (38)
where $\lambda$ is a constant. [^7] In the purely Riemannian setting ($T^{\sigma}{}_{\mu\nu}=0=Q_{\mu\nu\sigma}$), (38) implies $\lambda=\tfrac{2\Lambda}{n-2}$ in $n$ dimensions if the field equations are $R_{\mu\nu}-\tfrac{1}{2}Rg_{\mu\nu}+\Lambda g_{\mu\nu}=0$. In the present metric-affine setting, we simply take (38) as a geometric condition. Using (38), the Ricci terms in (35) simplify as
$R^{k}{}_{\ \rho}\,R_{\gamma\lambda\sigma k}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}=\lambda\big(g^{k}{}_{\rho}R_{\gamma\lambda\sigma k}-g^{k}{}_{\sigma}R_{\gamma\lambda\rho k}\big)=\lambda\big(R_{\gamma\lambda\sigma\rho}-R_{\gamma\lambda\rho\sigma}\big).$ (39)
Since the Riemann tensor remains antisymmetric in its last two indices, $R_{\gamma\lambda\sigma\rho}=-R_{\gamma\lambda\rho\sigma}$, the right-hand side of (39) becomes
$\lambda\big(R_{\gamma\lambda\sigma\rho}-R_{\gamma\lambda\rho\sigma}\big)=2\lambda\,R_{\gamma\lambda\sigma\rho}.$ (40)
Substituting (40) into the general wave equation (35) yields the Einstein-space form:
$\Box R_{\gamma\lambda\rho\sigma}+\nabla_{\rho}\nabla^{\nu}R_{\gamma\lambda\sigma\nu}-\nabla_{\sigma}\nabla^{\nu}R_{\gamma\lambda\rho\nu}+\nabla^{\nu}\!\left(X_{\gamma\lambda\nu\rho\sigma}+Y_{\gamma\lambda\nu\rho\sigma}\right)$
$\quad+T^{\nu k}{}_{\rho}\,\nabla_{\nu}R_{\gamma\lambda\sigma k}-T^{\nu k}{}_{\sigma}\,\nabla_{\nu}R_{\gamma\lambda\rho k}$
$\quad+R^{k}{}_{\gamma\rho}{}^{\nu}R_{k\lambda\sigma\nu}+R^{k}{}_{\lambda\rho}{}^{\nu}R_{\gamma k\sigma\nu}+R^{k}{}_{\sigma\rho}{}^{\nu}R_{\gamma\lambda k\nu}$
$\quad-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}+2\lambda\,R_{\gamma\lambda\sigma\rho}=0.$ (41)
Equation (41) is the specialization of the metric–affine curvature wave equation to Einstein spaces. In the torsion-free and metric-compatible limit, the $X$- and $Y$-terms vanish, and one recovers the familiar Riemannian curvature wave equation in an Einstein background.
V.2 Wave equation in the torsion- and nonmetricity-free case
We now consider the Riemannian (Levi-Civita) limit in which both torsion and nonmetricity vanish,
$T^{\mu}{}_{\nu\sigma}=0,\qquad Q_{\mu\nu\sigma}=0.$ (42)
In this case, the connection reduces to the Levi-Civita one, and correspondingly, the extra source terms introduced in (26)–(27) disappear:
$X_{\gamma\lambda\nu\rho\sigma}=0,\qquad Y_{\gamma\lambda\nu\rho\sigma}=0.$ (43)
The Riemann tensor then enjoys its standard Riemannian symmetries, namely,
$R_{\mu\nu\rho\sigma}=-R_{\nu\mu\rho\sigma}=-R_{\mu\nu\sigma\rho}=R_{\rho\sigma\mu\nu},$ (44)
together with the first Bianchi identity
$R^{k}{}_{\sigma\rho\nu}+R^{k}{}_{\rho\nu\sigma}+R^{k}{}_{\nu\sigma\rho}=0.$ (45)
Moreover, the contracted second Bianchi identity reduces to its familiar form,
$\nabla^{\gamma}R_{\gamma\lambda\rho\sigma}=\nabla_{\rho}R_{\lambda\sigma}-\nabla_{\sigma}R_{\lambda\rho}.$ (46)
Starting from the general wave-type equation (35) and imposing (42)–(43), we obtain
$\Box R_{\gamma\lambda\rho\sigma}+R^{k}{}_{\gamma\rho}{}^{\nu}R_{k\lambda\sigma\nu}+R^{k}{}_{\lambda\rho}{}^{\nu}R_{\gamma k\sigma\nu}+R^{k}{}_{\sigma\rho}{}^{\nu}R_{\gamma\lambda k\nu}+R^{k}{}_{\ \rho}\,R_{\gamma\lambda\sigma k}$
$\quad-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}$
$\quad+\nabla_{\rho}\nabla^{\nu}R_{\gamma\lambda\sigma\nu}-\nabla_{\sigma}\nabla^{\nu}R_{\gamma\lambda\rho\nu}=0.$ (47)
Using the exchange symmetry $R_{\rho\sigma\mu\nu}=R_{\mu\nu\rho\sigma}$ in (44) and the first Bianchi identity (45), the purely quadratic curvature terms in (47) can be brought to the standard symmetric form (see, e.g., the classic Lichnerowicz identity). After these rearrangements, (47) may be written as
$\Box R_{\gamma\lambda\rho\sigma}+2\,R_{k\lambda\sigma\nu}\,R^{k}{}_{\gamma\rho}{}^{\nu}+2\,R_{k\lambda\rho\nu}\,R^{k}{}_{\gamma\sigma}{}^{\nu}+R_{\gamma\lambda k\nu}\,R^{k\nu}{}_{\rho\sigma}$
$\quad+R^{k}{}_{\ \rho}\,R_{\gamma\lambda\sigma k}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}+\nabla_{\rho}\nabla^{\nu}R_{\gamma\lambda\sigma\nu}-\nabla_{\sigma}\nabla^{\nu}R_{\gamma\lambda\rho\nu}=0.$ (48)
Finally, the remaining divergence terms are simplified using the contracted Bianchi identity (46). In particular,
$\nabla^{\nu}R_{\gamma\lambda\sigma\nu}=\nabla_{\lambda}R_{\gamma\sigma}-\nabla_{\gamma}R_{\lambda\sigma},\qquad\nabla^{\nu}R_{\gamma\lambda\rho\nu}=\nabla_{\lambda}R_{\gamma\rho}-\nabla_{\gamma}R_{\lambda\rho},$ (49)
so that (48) becomes
$\Box R_{\gamma\lambda\rho\sigma}+2\,R_{k\lambda\sigma\nu}\,R^{k}{}_{\gamma\rho}{}^{\nu}+2\,R_{k\lambda\rho\nu}\,R^{k}{}_{\gamma\sigma}{}^{\nu}+R_{\gamma\lambda k\nu}\,R^{k\nu}{}_{\rho\sigma}$
$\quad+R^{k}{}_{\ \rho}\,R_{\gamma\lambda\sigma k}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}+\nabla_{\rho}\!\left(\nabla_{\lambda}R_{\gamma\sigma}-\nabla_{\gamma}R_{\lambda\sigma}\right)-\nabla_{\sigma}\!\left(\nabla_{\lambda}R_{\gamma\rho}-\nabla_{\gamma}R_{\lambda\rho}\right)=0.$ (50)
Equivalently, expanding the last line yields the fully explicit form
$\Box R_{\gamma\lambda\rho\sigma}+2\,R_{k\lambda\sigma\nu}\,R^{k}{}_{\gamma\rho}{}^{\nu}+2\,R_{k\lambda\rho\nu}\,R^{k}{}_{\gamma\sigma}{}^{\nu}+R_{\gamma\lambda k\nu}\,R^{k\nu}{}_{\rho\sigma}+R^{k}{}_{\ \rho}\,R_{\gamma\lambda\sigma k}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}$
$\quad+\nabla_{\rho}\nabla_{\lambda}R_{\gamma\sigma}-\nabla_{\rho}\nabla_{\gamma}R_{\lambda\sigma}-\nabla_{\sigma}\nabla_{\lambda}R_{\gamma\rho}+\nabla_{\sigma}\nabla_{\gamma}R_{\lambda\rho}=0.$ (51)
Equation (51) is the standard curvature wave equation in Riemannian geometry that we discussed in the introduction section and serves as a consistency check of the general metric-affine expression derived in the previous subsection.
V.3 Wave equation in the nonmetricity-free case
We next consider the metric-compatible case in which nonmetricity vanishes,
$Q_{\mu\nu\sigma}=0,$ (52)
while torsion is allowed to be nonzero,
$T^{\mu}{}_{\nu\sigma}\neq 0.$ (53)
Since $Q_{\mu\nu\sigma}=0$, the connection is metric compatible, and the covariant derivative preserves the metric, $\nabla_{\mu}g_{\nu\sigma}=0$. Consequently, indices may be raised and lowered freely with $g_{\mu\nu}$ without generating extra nonmetricity terms. In particular, the nonmetricity-dependent tensor introduced in (27) vanishes identically:
$Y_{\gamma\lambda\nu\rho\sigma}=0.$ (54)
Even though the torsion is present, the metric compatibility implies that the curvature with all indices lowered is antisymmetric in its first pair as well as in its second pair:
$R_{\mu\nu\rho\sigma}=-R_{\nu\mu\rho\sigma},\qquad R_{\mu\nu\rho\sigma}=-R_{\mu\nu\sigma\rho}.$ (55)
However, in general, one still does not have the pair-exchange symmetry
$R_{\mu\nu\rho\sigma}\neq R_{\rho\sigma\mu\nu},$ (56)
so the divergence of the Riemann tensor with respect to its last index cannot, in general, be rewritten solely in terms of the Ricci tensor as in the Levi–Civita case. This is one of the main structural differences between torsionful (metric-compatible) geometries and Riemannian geometry. Imposing (52) (and hence (54)) in the general curvature wave equation (35), we obtain
$\Box R_{\gamma\lambda\rho\sigma}+\nabla_{\rho}\nabla^{\nu}R_{\gamma\lambda\sigma\nu}-\nabla_{\sigma}\nabla^{\nu}R_{\gamma\lambda\rho\nu}+\nabla^{\nu}X_{\gamma\lambda\nu\rho\sigma}$
$\quad+T^{\nu k}{}_{\rho}\,\nabla_{\nu}R_{\gamma\lambda\sigma k}-T^{\nu k}{}_{\sigma}\,\nabla_{\nu}R_{\gamma\lambda\rho k}$
$\quad+R^{k}{}_{\gamma\rho}{}^{\nu}R_{k\lambda\sigma\nu}+R^{k}{}_{\lambda\rho}{}^{\nu}R_{\gamma k\sigma\nu}+R^{k}{}_{\sigma\rho}{}^{\nu}R_{\gamma\lambda k\nu}+R^{k}{}_{\ \rho}\,R_{\gamma\lambda\sigma k}$
$\quad-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}=0.$ (57)
Equation (57) is the metric-compatible, torsionful specialization of the general metric-affine curvature wave equation; it will be of particular interest when we further restrict to the Einstein-Cartan theory in the next subsection.  [^8] For electromagnetic waves propagating on a background with nonvanishing torsion, see [17].
V.4 Wave equation in the torsion-free case
We finally consider the torsion-free sector of metric-affine geometry, in which
$T^{\mu}{}_{\nu\sigma}=0,$ (58)
while the nonmetricity is allowed to be nonvanishing,
$Q_{\mu\nu\sigma}\neq 0.$ (59)
Since the torsion vanishes, the torsion-dependent tensor defined in (26) disappears identically,
$X_{\gamma\lambda\nu\rho\sigma}=0,$ (60)
whereas the nonmetricity-dependent tensor $Y_{\gamma\lambda\nu\rho\sigma}$ generally remains nonzero. In the absence of torsion, the curvature tensor regains antisymmetry in its last two indices,
$R^{\mu}{}_{\nu\rho\sigma}=-R^{\mu}{}_{\nu\sigma\rho}$, (61)
but, due to nonmetricity, neither antisymmetry in the first index pair nor pair-exchange symmetry needs to hold. In particular,
$R_{\mu\nu\rho\sigma}\neq-R_{\nu\mu\rho\sigma},\qquad R_{\mu\nu\rho\sigma}\neq R_{\rho\sigma\mu\nu}.$ (62)
As a consequence, raising and lowering indices with the metric produces additional terms proportional to $Q_{\mu\nu\sigma}$, and the divergences of the Riemann tensor cannot be simplified using the standard Riemannian identities.
Imposing (58) and (60) in the general curvature wave equation (35), we obtain
$\Box R_{\gamma\lambda\rho\sigma}+\nabla_{\rho}\nabla^{\nu}R_{\gamma\lambda\sigma\nu}-\nabla_{\sigma}\nabla^{\nu}R_{\gamma\lambda\rho\nu}+\nabla^{\nu}Y_{\gamma\lambda\nu\rho\sigma}$
$\quad+R^{k}{}_{\gamma\rho}{}^{\nu}R_{k\lambda\sigma\nu}+R^{k}{}_{\lambda\rho}{}^{\nu}R_{\gamma k\sigma\nu}+R^{k}{}_{\sigma\rho}{}^{\nu}R_{\gamma\lambda k\nu}+R^{k}{}_{\ \rho}\,R_{\gamma\lambda\sigma k}$
$\quad-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}=0.$ (63)
Equation (63) is the torsion-free, nonmetricity including the curvature wave equation. Compared to the metric-compatible case (57), the only additional source term is the divergence of $Y_{\gamma\lambda\nu\rho\sigma}$, which encodes the effect of nonmetricity on the propagation of curvature. In the simultaneous limit $T^{\mu}{}_{\nu\sigma}=0$ and $Q_{\mu\nu\sigma}=0$, this equation reduces smoothly to the standard Riemannian curvature wave equation derived in Section V.2. In addition to the given constraints, namely, the vanishing of torsion and the nonvanishing of nonmetricity, one may further assume the vanishing of the Riemann tensor, which leads to Symmetric Teleparallel Gravity (STGR). Moreover, although general relativity has no local degrees of freedom in three dimensions, STGR does possess them [18].
VI CONCLUSIONS AND FURTHER DISCUSSIONS
In Riemannian geometry [6] equipped with the Levi-Civita connection [7], the curvature of spacetime is not only a measure of gravitational interaction but also a dynamical geometric object. As is well known, the Riemann tensor satisfies a covariant wave equation whose source consists of terms quadratic in curvature [1]. This equation, which follows from the Bianchi identities and the Ricci identity, provides a precise geometric statement of the fact that curvature propagates: in this sense, Riemannian geometry is already wavy.
In this work, we have extended this familiar picture to the most general metric-affine setting, where the affine connection possesses torsion and nonmetricity, in addition to curvature. Within this framework, we derived a wave-type equation satisfied by the Riemann tensor for a generic connection. The resulting equation exhibits a rich structure: besides the usual curvature–curvature interaction terms known from the Riemannian case, it contains additional contributions driven by the torsion and nonmetricity. These appear both as algebraic couplings to the curvature and as derivative (transport-like) terms, reflecting the fact that torsion and nonmetricity modify the way geometric information propagates across spacetime.
By analyzing several important limiting cases, we clarified how different geometric structures affect the propagation of curvature. In the purely Riemannian limit, where both the torsion and nonmetricity vanish, our general expression reduces smoothly to the standard curvature wave equation. In Einstein spaces, the equation simplifies further and acquires an effective mass-like term proportional to the cosmological constant. In metric-compatible but torsionful geometries [10], such as Einstein-Cartan theory [8], torsion acts as an additional source and transport mechanism for curvature waves. Conversely, in torsion-free but nonmetric geometries [12], the influence of nonmetricity enters through explicit derivative couplings that have no analog in Riemannian geometry.
Taken together, these results reinforce a unifying geometric viewpoint: Gravity can be described equivalently in terms of curvature, torsion, or nonmetricity [10], but once these structures are present, geometry itself becomes dynamic and wavy. Curvature does not merely encode static geometric information; it propagates, interacts, and is influenced by the full affine structure of spacetime.
Several directions for further investigation naturally emerge from this work. First, it would be interesting to study the physical interpretation of the additional torsion- and nonmetricity-induced terms in specific models of metric-affine gravity, particularly in relation to spin, dilation, and shear currents of matter. One can find works in the literature along this direction [19, 20] Second, linearizing the general curvature wave equation around the appropriate backgrounds may shed light on the propagation and polarization content of the geometric waves beyond general relativity. Finally, the role of these curvature wave equations in exact solutions, integrability properties, and conservation laws deserves further exploration.
We hope that the results presented here contribute to a deeper understanding of the dynamical aspects of geometry in gravitational theories and provide a useful bridge between the classical Riemannian picture and its most general metric-affine extensions.
APPENDIX A: DECOMPOSITION OF A GENERIC AFFINE CONNECTION
In this appendix, we derive the standard decomposition of a generic affine connection into its Levi-Civita, torsional, and nonmetric contributions. We assume that the spacetime metric $g_{\mu\nu}$ and the affine connection $\Gamma^{\lambda}{}_{\mu\nu}$ are independent.
A.1 The nonmetricity and the torsion
The nonmetricity tensor is defined by
$Q_{\sigma\mu\nu}:=\nabla_{\sigma}g_{\mu\nu}=\partial_{\sigma}g_{\mu\nu}-\Gamma^{\lambda}{}_{\mu\sigma}g_{\lambda\nu}-\Gamma^{\lambda}{}_{\nu\sigma}g_{\mu\lambda}.$ (64)
Cyclic permutations of the indices yield
$Q_{\mu\nu\sigma}$ $=\partial_{\mu}g_{\nu\sigma}-\Gamma^{\lambda}{}_{\nu\mu}g_{\lambda\sigma}-\Gamma^{\lambda}{}_{\sigma\mu}g_{\nu\lambda},$ (65)
$Q_{\nu\sigma\mu}$ $=\partial_{\nu}g_{\sigma\mu}-\Gamma^{\lambda}{}_{\sigma\nu}g_{\lambda\mu}-\Gamma^{\lambda}{}_{\mu\nu}g_{\sigma\lambda}.$ (66)
Combining (64)–(66), we obtain
$Q_{\sigma\mu\nu}+Q_{\mu\nu\sigma}-Q_{\nu\sigma\mu}={}$ $\partial_{\sigma}g_{\mu\nu}+\partial_{\mu}g_{\nu\sigma}-\partial_{\nu}g_{\sigma\mu}$
$-\Gamma^{\lambda}{}_{\mu\sigma}g_{\lambda\nu}-\Gamma^{\lambda}{}_{\nu\sigma}g_{\mu\lambda}-\Gamma^{\lambda}{}_{\nu\mu}g_{\lambda\sigma}$
$-\Gamma^{\lambda}{}_{\sigma\mu}g_{\nu\lambda}+\Gamma^{\lambda}{}_{\sigma\nu}g_{\lambda\mu}+\Gamma^{\lambda}{}_{\mu\nu}g_{\sigma\lambda}.$ (67)
Assuming a symmetric metric $g_{\mu\nu}=g_{\nu\mu}$, this simplifies to
$Q_{\sigma\mu\nu}+Q_{\mu\nu\sigma}-Q_{\nu\sigma\mu}={}$ $\partial_{\sigma}g_{\mu\nu}+\partial_{\mu}g_{\nu\sigma}-\partial_{\nu}g_{\sigma\mu}$
$-\big(\Gamma^{\lambda}{}_{\mu\sigma}+\Gamma^{\lambda}{}_{\sigma\mu}\big)g_{\nu\lambda}+\big(\Gamma^{\lambda}{}_{\sigma\nu}-\Gamma^{\lambda}{}_{\nu\sigma}\big)g_{\mu\lambda}$
$+\big(\Gamma^{\lambda}{}_{\mu\nu}-\Gamma^{\lambda}{}_{\nu\mu}\big)g_{\sigma\lambda}.$ (68)
The torsion tensor is defined as
$T^{\lambda}{}_{\mu\nu}:=2\Gamma^{\lambda}{}_{[\mu\nu]}=\Gamma^{\lambda}{}_{\mu\nu}-\Gamma^{\lambda}{}_{\nu\mu}$. (69)
Using (69), equation (68) becomes
$Q_{\sigma\mu\nu}+Q_{\mu\nu\sigma}-Q_{\nu\sigma\mu}={}$ $\partial_{\sigma}g_{\mu\nu}+\partial_{\mu}g_{\nu\sigma}-\partial_{\nu}g_{\sigma\mu}$
$-2\Gamma^{\lambda}{}_{\mu\sigma}g_{\nu\lambda}-T_{\nu\sigma\mu}+T_{\mu\sigma\nu}+T_{\sigma\mu\nu}.$ (70)
A.2 Extraction of the affine connection
Contracting (70) with $\tfrac{1}{2}g^{\nu\beta}$ yields
[…]
The first three terms on the right-hand side form the Levi-Civita connection,
$^{c}\Gamma^{\beta}{}_{\mu\sigma}:=\frac{1}{2}g^{\nu\beta}\big(\partial_{\sigma}g_{\mu\nu}+\partial_{\mu}g_{\nu\sigma}-\partial_{\nu}g_{\sigma\mu}\big),$ (72)
which is symmetric in its lower indices. Using the symmetries
$Q_{\mu\nu\sigma}=Q_{\mu\sigma\nu},\qquad T_{\mu\nu\sigma}=-T_{\mu\sigma\nu},$ (73)
equation (71) can be solved for the full affine connection:
[…]
A.3 The disformation and contorsion
It is convenient to introduce the disformation tensor
[…]
and the contorsion tensor
[…]
With these definitions, the generic affine connection decomposes as
$\boxed{\Gamma^{\beta}{}_{\mu\sigma}={}^{c}\Gamma^{\beta}{}_{\mu\sigma}+L^{\beta}{}_{\mu\sigma}+K^{\beta}{}_{\mu\sigma}.}$ (77)
Equation (77) expresses the most general affine connection as the sum of the Levi–Civita connection, a purely nonmetric contribution, and a purely torsional contribution.
A.4 Properties of the torsion, non-metricity, contorsion, and the disformation tensors
  • The Non-metricity tensor is symmetric in its last two indices:
    $Q_{\sigma\mu\nu}=Q_{\sigma\nu\mu}.$ (78)
  • The disformation tensor is symmetric in its last two indices:
    $L_{\sigma\mu\nu}=L_{\sigma\nu\mu}.$ (79)
PROOF: By definition, one has
$L_{\sigma\mu\nu}$ $=\frac{1}{2}\left(Q_{\sigma\mu\nu}-Q_{\mu\sigma\nu}-Q_{\nu\sigma\mu}\right)$ (80)
$=\frac{1}{2}\left(Q_{\sigma\nu\mu}-Q_{\nu\sigma\mu}-Q_{\mu\sigma\nu}\right)$
$=L_{\sigma\nu\mu}.$
  • contorsion is antisymmetric in its first two indices:
    $K_{\sigma\mu\nu}=-K_{\mu\sigma\nu}.$ (81)
PROOF: By definition, one has
$K_{\sigma\mu\nu}$ $=\frac{1}{2}\left(T_{\sigma\mu\nu}-T_{\mu\sigma\nu}-T_{\nu\sigma\mu}\right)$ (82)
$=\frac{1}{2}\left(-T_{\sigma\nu\mu}+T_{\mu\nu\sigma}+T_{\nu\mu\sigma}\right),$
and equivalently, it becomes
$K_{\sigma\mu\nu}$ $=\frac{1}{2}\left(-T_{\mu\sigma\nu}+T_{\sigma\mu\nu}+T_{\nu\mu\sigma}\right)$ (83)
$=-\frac{1}{2}\left(T_{\mu\sigma\nu}-T_{\sigma\mu\nu}-T_{\nu\mu\sigma}\right)$
$=-K_{\mu\sigma\nu}.$
  • Difference of the contorsion tensors:
$K_{\mu\sigma\nu}-K_{\nu\sigma\mu}=-T_{\sigma\mu\nu}=T_{\sigma\nu\mu}.$ (84)
PROOF: The left-hand side reads
$K_{\mu\sigma\nu}-K_{\nu\sigma\mu}$ $=\frac{1}{2}\left(T_{\mu\sigma\nu}-T_{\sigma\mu\nu}-T_{\nu\mu\sigma}-T_{\nu\sigma\mu}+T_{\sigma\nu\mu}+T_{\mu\nu\sigma}\right)$ (85)
$=\frac{1}{2}\left(T_{\mu\sigma\nu}-T_{\sigma\mu\nu}-T_{\nu\mu\sigma}+T_{\nu\mu\sigma}-T_{\sigma\mu\nu}-T_{\mu\sigma\nu}\right)$
$=\frac{1}{2}\left(-T_{\sigma\mu\nu}-T_{\sigma\mu\nu}\right)$
$=-T_{\sigma\mu\nu}.$
  • One more difference of the contorsion tensors can be written as:
$K_{\mu\nu\sigma}-K_{\mu\sigma\nu}=T_{\mu\nu\sigma}.$ (86)
PROOF: The left-hand side reads
$K_{\mu\nu\sigma}-K_{\mu\sigma\nu}$ $=\frac{1}{2}\left(-T_{\mu\sigma\nu}+T_{\sigma\mu\nu}+T_{\nu\mu\sigma}+T_{\mu\nu\sigma}-T_{\nu\mu\sigma}-T_{\sigma\mu\nu}\right)$ (87)
$=\frac{1}{2}\left(-T_{\mu\sigma\nu}+T_{\mu\nu\sigma}\right)$
$=\frac{1}{2}\left(T_{\mu\nu\sigma}+T_{\mu\nu\sigma}\right)$
$=T_{\mu\nu\sigma}.$
  • Equality of the sums:
    $K_{\mu\sigma\nu}+K_{\sigma\nu\mu}+K_{\nu\mu\sigma}=\frac{1}{2}\left(T_{\mu\sigma\nu}+T_{\sigma\nu\mu}+T_{\nu\mu\sigma}\right).$ (88)
PROOF: By definition, the left-hand side reads
$K_{\mu\sigma\nu}+K_{\sigma\nu\mu}+K_{\nu\mu\sigma}=\frac{1}{2}$ $\left(T_{\mu\sigma\nu}-T_{\sigma\mu\nu}-T_{\nu\mu\sigma}+T_{\sigma\nu\mu}-T_{\nu\sigma\mu}\right$. (89)
$\left.-T_{\mu\sigma\nu}+T_{\nu\mu\sigma}-T_{\mu\nu\sigma}-T_{\sigma\nu\mu}\right)$
$=$ $\frac{1}{2}\left(-T_{\sigma\mu\nu}-T_{\nu\sigma\mu}-T_{\mu\nu\sigma}\right)$
$=$ $\frac{1}{2}\left(T_{\sigma\nu\mu}+T_{\nu\mu\sigma}+T_{\mu\sigma\nu}\right).$
Additionally, using the identity $K_{\mu\sigma\nu}-K_{\nu\sigma\mu}=T_{\sigma\nu\mu},$ we can express
$K_{\mu\sigma\nu}+K_{\sigma\nu\mu}+K_{\nu\mu\sigma}$ $=K_{\mu\sigma\nu}-K_{\nu\sigma\mu}+K_{\nu\mu\sigma}$ (90)
$=T_{\sigma\nu\mu}+K_{\nu\mu\sigma},$
which yields the following expression
$K_{\mu\sigma\nu}+K_{\sigma\nu\mu}+K_{\nu\mu\sigma}$ $=T_{\sigma\nu\mu}+\frac{1}{2}\left(T_{\nu\mu\sigma}-T_{\mu\nu\sigma}-T_{\sigma\nu\mu}\right)$ (91)
$=\frac{1}{2}\left(T_{\nu\mu\sigma}+T_{\sigma\nu\mu}+T_{\mu\sigma\nu}\right).$
APPENDIX B: THE CURVATURE TENSORS
B.1 The Riemann tensor
Given a generic affine connection $\Gamma^{\rho}{}_{\mu\nu}$, the associated Riemann curvature tensor [6] is defined by
$R^{\rho}{}_{\mu\sigma\nu}:=\partial_{\sigma}\Gamma^{\rho}{}_{\mu\nu}-\partial_{\nu}\Gamma^{\rho}{}_{\mu\sigma}+\Gamma^{\rho}{}_{\lambda\sigma}\Gamma^{\lambda}{}_{\mu\nu}-\Gamma^{\rho}{}_{\lambda\nu}\Gamma^{\lambda}{}_{\mu\sigma}$. (92)
Throughout, we use the decomposition (77) derived in Appendix A. Substituting (77) into (92) and expanding, one may organize the result as the Levi-Civita curvature, along with the terms involving $K$ and $L$. The Levi-Civita Riemann tensor is
$^{c}R^{\rho}{}_{\mu\sigma\nu}:=\partial_{\sigma}\,{}^{c}\Gamma^{\rho}{}_{\mu\nu}-\partial_{\nu}\,{}^{c}\Gamma^{\rho}{}_{\mu\sigma}+{}^{c}\Gamma^{\rho}{}_{\lambda\sigma}\,{}^{c}\Gamma^{\lambda}{}_{\mu\nu}-{}^{c}\Gamma^{\rho}{}_{\lambda\nu}\,{}^{c}\Gamma^{\lambda}{}_{\mu\sigma}$. (93)
For compactness, it is convenient to introduce the distortion tensor
$N^{\rho}{}_{\mu\nu}:=K^{\rho}{}_{\mu\nu}+L^{\rho}{}_{\mu\nu}$, (94)
which is a $(1,2)$ tensor. Then the curvature decomposition takes the standard form
$R^{\rho}{}_{\mu\sigma\nu}={}^{c}R^{\rho}{}_{\mu\sigma\nu}+{}^{c}\nabla_{\sigma}N^{\rho}{}_{\mu\nu}-{}^{c}\nabla_{\nu}N^{\rho}{}_{\mu\sigma}+N^{\rho}{}_{\lambda\sigma}N^{\lambda}{}_{\mu\nu}-N^{\rho}{}_{\lambda\nu}N^{\lambda}{}_{\mu\sigma}$, (95)
where $^{c}\nabla$ denotes the Levi–Civita covariant derivative. Expanding $N^{\rho}{}_{\mu\nu}=K^{\rho}{}_{\mu\nu}+L^{\rho}{}_{\mu\nu}$ in (95), we obtain an explicit expression in terms of contorsion and disformation:
[…]
It is sometimes useful to rewrite the mixed torsion contribution using the identity
$T^{\lambda}{}_{\nu\sigma}=\Gamma^{\lambda}{}_{\nu\sigma}-\Gamma^{\lambda}{}_{\sigma\nu}=K^{\lambda}{}_{\nu\sigma}-K^{\lambda}{}_{\sigma\nu}$, (97)
which follows from the definition of contorsion. With this, one may equivalently present the decomposition in the form used in the main text:
[…]
where $\nabla$ is the covariant derivative built from the full connection $\Gamma^{\rho}{}_{\mu\nu}$. The two representations (96) and (98) are equivalent; the former is often cleaner for algebraic manipulations (it uses $^{c}\nabla$), while the latter is convenient for keeping expressions manifestly covariant with respect to the full connection.
Finally, irrespective of torsion and nonmetricity, the curvature defined by (92) remains antisymmetric in its last two indices:
$R^{\rho}{}_{\mu\sigma\nu}=-R^{\rho}{}_{\mu\nu\sigma}$. (99)
A quick check of curvature symmetries
In the metric–affine case, we obtained the Riemann tensor in the form
[…]
By definition (92), the curvature is always antisymmetric in its last two indices:
$R^{\rho}{}_{\mu\nu\sigma}=-R^{\rho}{}_{\mu\sigma\nu}$. (101)
The symmetry under the exchange of the first index pair is subtler. In the most general metric–affine case ($Q_{\alpha\mu\nu}\neq 0$), lowering indices involves the full connection and does not commute with covariant differentiation; consequently, $R_{\mu\nu\rho\sigma}$ need not be antisymmetric in $\mu,\nu$. More explicitly, one generally has
$R_{\mu\nu\rho\sigma}\neq-R_{\nu\mu\rho\sigma},$ (102)
and the pair-exchange symmetry also fails in general,
$R_{\mu\nu\rho\sigma}\neq R_{\rho\sigma\mu\nu}.$ (103)
On the other hand, in the metric-compatible case ($Q_{\alpha\mu\nu}=0$), one can lower the indices without generating extra nonmetricity terms. Then the curvature with all indices lowered regains antisymmetry in the first pair,
$Q_{\alpha\mu\nu}=0\qquad\Longrightarrow\qquad R_{\mu\nu\rho\sigma}=-R_{\nu\mu\rho\sigma},$ (104)
while the pair-exchange symmetry $R_{\mu\nu\rho\sigma}=R_{\rho\sigma\mu\nu}$ may still fail when torsion is present.
B2. The Ricci tensor
The Ricci tensor is defined by the contraction of the first and third indices,
$R_{\mu\nu}:=R^{\sigma}{}_{\mu\sigma\nu},\qquad{}^{c}R_{\mu\nu}:={}^{c}R^{\sigma}{}_{\mu\sigma\nu}$. (105)
Contracting (98) accordingly, one obtains
[…]
Here we used that the disformation tensor is symmetric in its last two indices, $L^{\rho}{}_{\mu\nu}=L^{\rho}{}_{\nu\mu}$, while the contorsion satisfies $K_{\rho\mu\nu}=-K_{\mu\rho\nu}$.
B3. The scalar curvature
The scalar curvature is $R=g^{\mu\nu}R_{\mu\nu}$. Writing also $^{c}R=g^{\mu\nu}{}^{c}R_{\mu\nu}$, we find from (106) that:
[…]
$\quad+K^{\lambda\mu\sigma}K_{\sigma\mu\lambda}+L^{\lambda\mu\sigma}L_{\sigma\mu\lambda}+K^{\mu\lambda}{}_{\mu}K^{\sigma}{}_{\lambda\sigma}+K^{\mu\lambda}{}_{\mu}L^{\sigma}{}_{\lambda\sigma}-L^{\lambda\mu}{}_{\mu}K^{\sigma}{}_{\lambda\sigma}-L^{\lambda\mu}{}_{\mu}L^{\sigma}{}_{\lambda\sigma}$. (107)
Equation (107) is the scalar curvature in terms of Levi–Civita curvature, plus the torsion and the nonmetricity contributions. Note that in the generic case one has $\nabla_{\sigma}g^{\mu\nu}\neq 0$; hence, the divergence terms in (107) should be kept as written.
Metric-compatible limit
When the connection is metric compatible ($Q_{\alpha\mu\nu}=0$), the disformation vanishes, $L^{\rho}{}_{\mu\nu}=0$, and (107) reduces to
$R={}$ $^{c}R+g^{\mu\nu}\nabla_{\sigma}K^{\sigma}{}_{\mu\nu}-\nabla_{\mu}K^{\sigma}{}_{\mu\sigma}+K^{\lambda\mu\sigma}K_{\sigma\mu\lambda}+K^{\mu\lambda}{}_{\mu}K^{\sigma}{}_{\lambda\sigma}$. (108)
Using $Q_{\alpha\mu\nu}=0$ one may move the metric through the covariant derivative, so $g^{\mu\nu}\nabla_{\sigma}K^{\sigma}{}_{\mu\nu}=\nabla_{\sigma}K^{\sigma\mu}{}_{\mu}$, and therefore
$R=^{c}R+2\,\nabla_{\sigma}\!K^{\sigma\mu}{}_{\mu}+K^{\lambda\mu\sigma}K_{\sigma\mu\lambda}+K^{\mu\lambda}{}_{\mu}K^{\sigma}{}_{\lambda\sigma}$. (109)
APPENDIX C: THE RICCI IDENTITY FOR THE COVARIANT DERIVATIVES
In this appendix, we derive the commutator (Ricci) identity for a generic affine connection with torsion and possibly nonmetricity. The nonmetricity affects how the indices are raised and lowered, but the commutator formula itself is determined by the torsion and curvature.
C.1 Warm-up: action on a scalar
Let $f$ be a scalar field. Since $\nabla_{\mu}f=\partial_{\mu}f$, we compute
$[\nabla_{\mu},\nabla_{\nu}]f$ $=\nabla_{\mu}\nabla_{\nu}f-\nabla_{\nu}\nabla_{\mu}f$
$=\nabla_{\mu}(\partial_{\nu}f)-\nabla_{\nu}(\partial_{\mu}f)$
$=\partial_{\mu}\partial_{\nu}f-\Gamma^{\lambda}{}_{\nu\mu}\partial_{\lambda}f-\partial_{\nu}\partial_{\mu}f+\Gamma^{\lambda}{}_{\mu\nu}\partial_{\lambda}f$
$=\left(\Gamma^{\lambda}{}_{\mu\nu}-\Gamma^{\lambda}{}_{\nu\mu}\right)\partial_{\lambda}f=T^{\lambda}{}_{\mu\nu}\,\nabla_{\lambda}f,$ (110)
where the torsion tensor is $T^{\lambda}{}_{\mu\nu}:=\Gamma^{\lambda}{}_{\mu\nu}-\Gamma^{\lambda}{}_{\nu\mu}$. Thus, unlike the Levi-Civita case, the commutator does not vanish even for scalars.
C.2 Action on a covector and a vector
For a covector $V_{\sigma}$ one finds
$\nabla_{\mu},\nabla_{\nu}]V_{\sigma}=T^{\lambda}{}_{\mu\nu}\,\nabla_{\lambda}V_{\sigma}+R^{\lambda}{}_{\sigma\nu\mu}\,V_{\lambda},$ (111)
where the Riemann tensor $R^{\lambda}{}_{\sigma\nu\mu}$ is defined by (92). For a vector $V^{\sigma}$ one similarly obtains
$\nabla_{\mu},\nabla_{\nu}]V^{\sigma}=T^{\lambda}{}_{\mu\nu}\,\nabla_{\lambda}V^{\sigma}-R^{\sigma}{}_{\lambda\nu\mu}\,V^{\lambda}.$ (112)
The sign difference between (111) and (112) is the usual one: curvature acts with the opposite sign on the covariant versus contravariant indices.
C.3 General Ricci identity
For a general tensor field $A^{\sigma\rho\cdots}{}_{\lambda k\cdots}$, repeated application of the Leibniz rule together with the results of Subsections C.1 and C.2 yields the Ricci (commutator) identity for a generic affine connection with torsion:
$[\nabla_{\mu},\nabla_{\nu}]\,A^{\sigma\rho\cdots}{}_{\lambda k\cdots}$={} $T^{\gamma}{}_{\mu\nu}\,\nabla_{\gamma}A^{\sigma\rho\cdots}{}_{\lambda k\cdots}$
$-R^{\sigma}{}_{\gamma\nu\mu}\,A^{\gamma\rho\cdots}{}_{\lambda k\cdots}-R^{\rho}{}_{\gamma\nu\mu}\,A^{\sigma\gamma\cdots}{}_{\lambda k\cdots}$-$\cdots$
$+R^{\gamma}{}_{\lambda\nu\mu}\,A^{\sigma\rho\cdots}{}_{\gamma k\cdots}+R^{\gamma}{}_{k\nu\mu}\,A^{\sigma\rho\cdots}{}_{\lambda\gamma\cdots}$$\cdots$. (113)
Here, each contravariant index contributes a curvature term with a minus sign, while each covariant index contributes a curvature term with a plus sign; the ellipsis indicates the obvious continuation to all remaining indices of the tensor.
Equation (113) is the form used throughout the main text. In the torsion-free limit, $T^{\gamma}{}_{\mu\nu}=0$, it reduces to the standard Ricci identity of Riemannian geometry.
APPENDIX D: BIANCHI IDENTITIES
In this appendix, we derive the generalized Bianchi identities for a generic affine connection with torsion. The derivation is based on the Jacobi identity satisfied by covariant derivatives.
D.1 Jacobi identity
For any vector field $V^{\mu}$, the covariant derivatives satisfy
$\left[\left[\nabla_{\sigma},\nabla_{\rho}\right],\nabla_{\nu}\right]V^{\mu}+\left[\left[\nabla_{\rho},\nabla_{\nu}\right],\nabla_{\sigma}\right]V^{\mu}+\left[\left[\nabla_{\nu},\nabla_{\sigma}\right],\nabla_{\rho}\right]V^{\mu}=0.$ (114)
This identity will lead simultaneously to the first and second Bianchi identities.
D.2 Evaluation of the nested commutators
We begin by evaluating the first term in (114):
$\left[\left[\nabla_{\sigma},\nabla_{\rho}\right],\nabla_{\nu}\right]V^{\mu}=\left[\nabla_{\sigma},\nabla_{\rho}\right]\nabla_{\nu}V^{\mu}-\nabla_{\nu}\left[\nabla_{\sigma},\nabla_{\rho}\right]V^{\mu}.$ (115)
Using the Ricci identity derived in Appendix C, one finds
$\left[\left[\nabla_{\sigma},\nabla_{\rho}\right],\nabla_{\nu}\right]V^{\mu}={}$ $T^{\lambda}{}_{\sigma\rho}\nabla_{\lambda}\nabla_{\nu}V^{\mu}+R^{\lambda}{}_{\nu\rho\sigma}\nabla_{\lambda}V^{\mu}-R^{\mu}{}_{\lambda\rho\sigma}\nabla_{\nu}V^{\lambda}$ (116)
$-\nabla_{\nu}\!\left(T^{\lambda}{}_{\sigma\rho}\nabla_{\lambda}V^{\mu}-R^{\mu}{}_{\lambda\rho\sigma}V^{\lambda}\right).$
Expanding the last term and regrouping yields
$\left[\left[\nabla_{\sigma},\nabla_{\rho}\right],\nabla_{\nu}\right]V^{\mu}={}$ $T^{\lambda}{}_{\sigma\rho}\left[\nabla_{\lambda},\nabla_{\nu}\right]V^{\mu}+\left(R^{\lambda}{}_{\nu\rho\sigma}-\nabla_{\nu}T^{\lambda}{}_{\sigma\rho}\right)\nabla_{\lambda}V^{\mu}$ (117)
$+V^{\lambda}\nabla_{\nu}R^{\mu}{}_{\lambda\rho\sigma}$.
Applying the Ricci identity once more, we arrive at
$\left[\left[\nabla_{\sigma},\nabla_{\rho}\right],\nabla_{\nu}\right]V^{\mu}={}$ $\left(R^{\lambda}{}_{\nu\rho\sigma}-\nabla_{\nu}T^{\lambda}{}_{\sigma\rho}+T^{\gamma}{}_{\sigma\rho}T^{\lambda}{}_{\gamma\nu}\right)\nabla_{\lambda}V^{\mu}$ (118)
$+\left(\nabla_{\nu}R^{\mu}{}_{\lambda\rho\sigma}-R^{\mu}{}_{\lambda\nu\gamma}T^{\gamma}{}_{\sigma\rho}\right)V^{\lambda}.$
The remaining two nested commutators are obtained by cyclic permutations of $(\sigma,\rho,\nu)$.
D.3 Generalized Bianchi identities
Substituting the three cyclic permutations of (118) into the Jacobi identity (114) and collecting terms proportional to $\nabla_{\lambda}V^{\mu}$ and $V^{\lambda}$ separately, one obtains
$R^{\lambda}{}_{\nu\rho\sigma}+R^{\lambda}{}_{\sigma\nu\rho}+R^{\lambda}{}_{\rho\sigma\nu}-\nabla_{\nu}T^{\lambda}{}_{\sigma\rho}-\nabla_{\sigma}T^{\lambda}{}_{\rho\nu}-\nabla_{\rho}T^{\lambda}{}_{\nu\sigma}$
$\qquad+T^{\gamma}{}_{\sigma\rho}T^{\lambda}{}_{\gamma\nu}+T^{\gamma}{}_{\rho\nu}T^{\lambda}{}_{\gamma\sigma}+T^{\gamma}{}_{\nu\sigma}T^{\lambda}{}_{\gamma\rho}=0,$ (119)
and
$\nabla_{\nu}R^{\mu}{}_{\lambda\rho\sigma}+\nabla_{\rho}R^{\mu}{}_{\lambda\sigma\nu}+\nabla_{\sigma}R^{\mu}{}_{\lambda\nu\rho}$
$\qquad-R^{\mu}{}_{\lambda\nu\gamma}T^{\gamma}{}_{\sigma\rho}-R^{\mu}{}_{\lambda\sigma\gamma}T^{\gamma}{}_{\rho\nu}-R^{\mu}{}_{\lambda\rho\gamma}T^{\gamma}{}_{\nu\sigma}=0.$ (120)
D.4 Compact antisymmetric form
Using antisymmetrization brackets, the first Bianchi identity (119) can be written compactly as
$R^{\lambda}{}_{[\nu\rho\sigma]}=-\nabla_{[\nu}T^{\lambda}{}_{\rho\sigma]}+T^{\gamma}{}_{[\nu\rho}T^{\lambda}{}_{\sigma]\gamma}$. (121)
Similarly, the second Bianchi identity (120) takes the compact form
$\nabla_{[\nu}R^{\mu}{}_{|\lambda|\rho\sigma]}=R^{\mu}{}_{\lambda\gamma[\nu}T^{\gamma}{}_{\rho\sigma]}$. (122)
In the torsion-free limit $T^{\lambda}{}_{\mu\nu}=0$, these identities reduce to the familiar Bianchi identities of Riemannian geometry.
APPENDIX E: TELEPARALLEL GRAVITY
Recall the generic expression for the Riemann tensor
$R_{\ \mu\sigma\nu}^{\rho}=$ ${}^{c}R_{\ \mu\sigma\nu}^{\rho}+\nabla_{\sigma}K_{\ \mu\nu}^{\rho}+\nabla_{\sigma}L_{\ \mu\nu}^{\rho}-\nabla_{\nu}K_{\ \mu\sigma}^{\rho}-\nabla_{\nu}L^{\rho}{}_{\mu\sigma}$ (123)
[…]
Using the splitting of the generic connection (77) and introducing the distortion tensor $D^{\mu}{}_{\nu\sigma}$
$D^{\mu}{}_{\nu\sigma}:=\Gamma^{\mu}{}_{\nu\sigma}-^{c}\Gamma^{\mu}{}_{\nu\sigma}=K^{\mu}{}_{\nu\sigma}+L^{\mu}{}_{\nu\sigma}$, (124)
we can rewrite the Riemann tensor as follows
[…]
$+D^{\lambda}{}_{\mu\sigma}D^{\rho}{}_{\lambda\nu}-D^{\lambda}{}_{\mu\nu}D^{\rho}{}_{\lambda\sigma}$.
Or equivalently, one arrives at
[…]
where one can use the symmetry $L^{\lambda}{}_{v\sigma}=L_{\ \sigma v}^{\lambda}$ in the last equation. Then the last equation reduces to
$R_{\ \mu\sigma\nu}^{\rho}=$ $^{c}R_{\ \mu\sigma\nu}^{\rho}+\nabla_{\sigma}D_{\ \mu\nu}^{\rho}-\nabla_{\nu}D^{\rho}{}_{\mu\sigma}$
[…]
E.1 Metric teleparallel connection
Metric teleparallel gravity [10] is described via the following properties of the tensor fields:
  • $R^{\mu}{}_{\nu\rho\sigma}=0,$
  • $Q_{\mu\nu\rho}=0\mathbf{\ \ \ \ \ }\Rightarrow\mathbf{\ \ \ \ \ }L^{\mu}{}_{\rho\sigma}=0,$
  • $T^{\mu}{}_{\nu\sigma}\neq 0\mathbf{\ \ \ \ \ }\Rightarrow\mathbf{\ \ \ \ \ }K^{\mu}{}_{\nu\sigma}\neq 0.$
In this case, the contributions to the distortion tensor come from the Levi Civita connection and the torsion tensor. Particularly, one has the equality of the distortion and the contorsion. Namely
$D^{\mu}{}_{v\sigma}=K^{\mu}{}_{v\sigma}$. (128)
Since $R^{\mu}{}_{\nu\rho\sigma}=0$ and $D^{\mu}{}_{\nu\sigma}=K^{\mu}{}_{\nu\sigma}$ one has
$0=$ ${}^{c}R^{\rho}{}_{\mu\sigma\nu}+\nabla_{\sigma}K^{\rho}{}_{\mu\nu}-\nabla_{\nu}K^{\rho}{}_{\mu\nu}+K^{\lambda}{}_{\mu\sigma}K^{\rho}{}_{\lambda\nu}$ (129)
[…]
Therefore, the Riemann tensor of the Levi-Civita connection, in terms of the contorsion tensor, reads
$^{c}R^{\rho}{}_{\mu\sigma\nu}$ $=\nabla_{\nu}K_{\ \mu\sigma}^{\rho}-\nabla_{\sigma}K^{\rho}{}_{\mu\nu}+K^{\lambda}{}_{\mu\nu}K^{\rho}{}_{\lambda\sigma}$
[…]
By way of contraction, one obtains the Ricci tensor as
$^{c}R_{\mu\nu}=$ $\nabla_{\nu}K^{\sigma}{}_{\mu\sigma}-\nabla_{\sigma}K^{\sigma}{}_{\mu\nu}+K^{\lambda}{}_{\mu\nu}K^{\sigma}{}_{\lambda\sigma}$ (131)
[…]
The scalar curvature then becomes
${}^{c}R=$ $\nabla_{\mu}K^{\sigma\mu}{}_{\sigma}-\nabla_{\sigma}K^{\sigma}{}_{\mu}{}^{\mu}+K^{\lambda}{}_{\mu}{}^{\mu}K^{\sigma}{}_{\lambda\sigma}$ (132)
[…]
Remember the antisymmetry property of the K-tensor: $K_{\mu\sigma\nu}=-K_{\sigma\mu\nu}$. So, we can write
$\begin{gathered}{}^{c}R=2\nabla_{\mu}K^{\sigma\mu}{}_{\sigma}-K_{\mu}{}^{\lambda\mu}K^{\sigma}{}_{\lambda\sigma}-K_{\lambda\mu\sigma}K^{\sigma\lambda\mu}+K^{\sigma\mu\lambda}K_{\lambda\sigma\mu}-K^{\sigma\mu\lambda}K_{\lambda\mu\sigma},\end{gathered}$ (133)
or in a more compact form
$\begin{gathered}{}^{c}R=2\nabla_{\mu}K^{\sigma\mu}{}_{\sigma-}K_{\mu}{}^{\lambda\mu}K^{\sigma}{}_{\lambda\sigma}+K^{\sigma\mu\lambda}K_{\lambda\sigma\mu}-K_{\lambda\mu\sigma}\left(K^{\sigma\lambda\mu}+K^{\sigma\mu\lambda}\right),\end{gathered}$ (134)
where the last two terms automatically vanish because of the symmetries. Then
$^{c}R=2\nabla_{\mu}K^{\sigma\mu}{}_{\sigma}-K_{\mu}{}^{\lambda\mu}K^{\sigma}{}_{\lambda\sigma}+K^{\sigma\mu\lambda}K_{\lambda\sigma\mu}.$ (135)
Using the explicit form
[…]
one has $T^{\mu}{}_{\nu\sigma}=-T^{\mu}{}_{\sigma\nu}$, and so
[…]
which reduces to
$K_{\ \mu\sigma}^{\sigma}=T_{\ \mu\sigma}^{\sigma}.$ (138)
Also, one has the following contraction:
$K^{\sigma\mu\lambda}K_{\lambda\sigma\mu}=\frac{1}{4}$ $\left(T^{\sigma\mu\lambda}-T^{\mu\sigma\lambda}-T^{\lambda\sigma\mu}\right)\left(T_{\lambda\sigma\mu}-T_{\sigma\lambda\mu}-T_{\mu\lambda\sigma}\right)$ (139)
$=\frac{1}{4}$ $\biggl(T^{\sigma\mu\lambda}T_{\lambda\sigma\mu}-T^{\sigma\mu\lambda}T_{\sigma\lambda\mu}-T^{\sigma\mu\lambda}T_{\mu\lambda\sigma}$
$-T^{\mu\sigma\lambda}T_{\lambda\sigma\mu}+T^{\mu\sigma\lambda}T_{\sigma\lambda\mu}+T^{\mu\sigma\lambda}T_{\mu\lambda\sigma}$
$-T^{\lambda\sigma\mu}T_{\lambda\sigma\mu}+T^{\lambda\sigma\mu}\left(T_{\sigma\lambda\mu}+T_{\mu\lambda\sigma}\right)\biggr),$
where the last two terms vanish because of the symmetries. Then we have
$K^{\sigma\mu\lambda}K_{\lambda\sigma\mu}=\frac{1}{4}$ $\biggl(T_{\lambda\sigma\mu}(T^{\sigma\mu\lambda}-T^{\mu\sigma\lambda}-T^{\lambda\sigma\mu})$ (140)
$+T_{\sigma\lambda\mu}(T^{\mu\sigma\lambda}-T^{\sigma\mu\lambda})+T_{\mu\lambda\sigma}(T^{\mu\sigma\lambda}-T^{\sigma\mu\lambda})\biggr).$
By renaming the indices, it has become
$K^{\sigma\mu\lambda}K_{\lambda\sigma\mu}$ $=\frac{1}{4}T_{\lambda\sigma\mu}\left(T^{\sigma\mu\lambda}-T^{\mu\sigma\lambda}-T^{\lambda\sigma\mu}+T^{\mu\lambda\sigma}-T^{\lambda\mu\sigma}+T^{\lambda\mu\sigma}-T^{\mu\lambda\sigma}\right),$ (141)
which reduces to
$K^{\sigma\mu\lambda}K_{\lambda\sigma\mu}=\frac{1}{4}T_{\lambda\sigma\mu}\left(T^{\sigma\mu\lambda}-T^{\mu\sigma\lambda}-T^{\lambda\sigma\mu}\right).$ (142)
Using the antisymmetry of the torsion, one has
$K^{\sigma\mu\lambda}K_{\lambda\sigma\mu}=\frac{1}{4}\left(T_{\lambda\sigma\mu}T^{\sigma\mu\lambda}+T_{\lambda\mu\sigma}T^{\mu\sigma\lambda}-T_{\lambda\sigma\mu}T^{\lambda\sigma\mu}\right).$ (143)
Finally, one ends up with the following expression
$K^{\sigma\mu\lambda}K_{\lambda\sigma\mu}$ $=\frac{1}{4}T_{\lambda\sigma\mu}\left(T^{\sigma\mu\lambda}-T^{\mu\sigma\lambda}-T^{\lambda\sigma\mu}\right)$
$=\frac{1}{4}\left(2T^{\sigma\mu\lambda}-T^{\lambda\sigma\mu}\right)T_{\lambda\sigma\mu}.$ (144)
Then, the scalar curvature can be rewritten as
$^{c}R=2\nabla_{\mu}K^{\sigma\mu}{}_{\sigma}-K_{\mu}{}^{\lambda\mu}K^{\sigma}{}_{\lambda\sigma}+K^{\sigma\mu\lambda}K_{\lambda\sigma\mu}=2\nabla_{\mu}T^{\sigma\mu}{}_{\sigma}-T_{\mu}{}^{\lambda\mu}T^{\sigma}{}_{\lambda\sigma}+\frac{1}{4}T_{\lambda\sigma\mu}\left(2T^{\sigma\mu\lambda}-T^{\lambda\sigma\mu}\right).$ (145)
The last equation implies that the Einstein-Hilbert action
$S_{EH}=\frac{1}{2R}\int d^{4}x\sqrt{-g}{}^{c}R,$ (146)
can be rewritten in terms of the torsion tensor instead of the Christoffel connection.
E.2 Symmetric teleparallel connection
Symmetric teleparallel gravity [12] connection-satisfies the following relations:
  • $R^{\mu}{}_{\nu\rho\sigma}=0,$
  • $T^{\mu}{}_{\nu\rho}=0\mathbf{\ \ \ \ \ }\Rightarrow\mathbf{\ \ \ \ \ }K^{\mu}{}_{\nu\rho}=0,$
  • $Q_{\mu\nu\rho}\neq 0\mathbf{\ \ \ \ \ }\Rightarrow\mathbf{\ \ \ \ \ }L^{\mu}{}_{\nu\rho}\neq 0.$
In this case, the distortion is equal to the L-tensor, $D^{\mu}{}_{\nu\sigma}=L^{\mu}{}_{\nu\sigma}$. Therefore, the Riemann tensor of the Christoffel connection can be expressed as
$^{c}R^{\rho}{}_{\mu\sigma\nu}$= $\nabla_{\nu}L^{\rho}{}_{\mu\sigma}-\nabla_{\sigma}L^{\rho}{}_{\mu\nu}+L^{\lambda}{}_{\mu\nu}L^{\rho}{}_{\lambda\sigma}$ (147)
[…]
and it yields the Ricci tensor of the Levi Civita connection as
$^{c}R_{\mu\nu}=$ $\nabla_{\nu}L^{\sigma}{}_{\mu\sigma}-\nabla_{\sigma}L^{\sigma}{}_{\mu\nu}+L^{\lambda}{}_{\mu\nu}L^{\sigma}{}_{\lambda\sigma}$ (148)
[…]
Note that we should be careful in raising the indices because of the non-vanishing of the nonmetricity tensor. The connection is not metric compatible, which means
$g^{\mu\nu}\nabla_{\sigma}L_{\penalty 10000\ \mu\nu}^{\sigma}\neq\nabla_{\sigma}L^{\sigma\mu}{}_{\mu}$. (149)
The Ricci scalar becomes
${}^{c}R=\nabla^{\mu}$ $L^{\sigma}{}_{\mu\sigma}-g{}^{\mu\nu}\nabla_{\sigma}L^{\sigma}{}_{\mu\nu}+L^{\lambda\mu}{}_{\mu}L^{\sigma}{}_{\lambda\sigma}$ (150)
[…]
Remember that $L_{\mu\nu\sigma}$ is symmetric in its last two indices. Hence, we can write
${}^{c}R=\nabla^{\mu}$ $L_{\mu\sigma-}^{\sigma}g^{\mu\nu}\nabla_{\sigma}L_{\mu\nu}^{\sigma}+L^{\lambda\mu}{}_{\mu}L^{\sigma}{}_{\lambda\sigma}$ (151)
$-L_{\lambda\mu\sigma}L^{\sigma\lambda\mu}+L_{\sigma\mu\lambda}\left(L^{\lambda\sigma\mu}-L^{\lambda\sigma\mu}\right),$
which reduces to the following
$^{c}R=\nabla^{\mu}L^{\sigma}{}_{\mu\sigma}-g^{\mu\nu}\nabla_{\sigma}L^{\sigma}{}_{\mu\nu}+L^{\lambda\mu}{}_{\mu}L^{\sigma}{}_{\sigma\lambda}-L_{\lambda\mu\sigma}L^{\sigma\lambda\mu}.$ (152)
Using this result, one can express the scalar curvature of the Levi-Civita connection in terms of $L$, and thus the Einstein-Hilbert action. That formulation provides an alternative approach to general relativity.
APPENDIX F: TETRADS, THE SPIN CONNECTION, AND THE TELEPARALLEL GEOMETRY
This appendix provides the minimal tetrad and the spin–connection [9] framework needed to interpret teleparallel gravity [10] and to connect the metric–affine wave equations derived in the main text with the gauge-theoretic formulations of the spacetime geometry. No new dynamics are introduced here; the purpose is conceptual unification.
F.1 Tetrads and the spacetime metric
We introduce the tetrad (vierbein) $e^{A}{}_{\mu}$ and its inverse $E_{A}{}^{\mu}$, satisfying
$E_{A}{}^{\mu}e^{A}{}_{\nu}=\delta^{\mu}{}_{\nu},\qquad E_{A}{}^{\mu}e^{B}{}_{\mu}=\delta_{A}^{B}.$ (153)
The spacetime metric is expressed as
$g_{\mu\nu}=\eta_{AB}\,e^{A}{}_{\mu}e^{B}{}_{\nu}$, (154)
where $\eta_{AB}=\mathrm{diag}(1,-1,-1,-1)$ is the Minkowski metric. This establishes the equivalence between the metric and the tetrad descriptions of spacetime geometry.
F.2 Spin connection and the affine connection
The spin connection [9], $\omega^{A}{}_{B\mu}$, relates to the spacetime affine connection via
[…]
This follows from the tetrad postulate
$\nabla_{\nu}e^{A}{}_{\mu}=\partial_{\nu}e^{A}{}_{\mu}+\omega^{A}{}_{B\nu}e^{B}{}_{\mu}-\Gamma^{\rho}{}_{\mu\nu}e^{A}{}_{\rho}=0,$ (156)
which ensures compatibility between spacetime and Lorentz indices.
F.3 Torsion and the curvature in differential form language
The torsion and curvature two-forms are defined as
$T^{A}=de^{A}+\omega^{A}{}_{B}\wedge e^{B},\qquad R^{A}{}_{B}=d\omega^{A}{}_{B}+\omega^{A}{}_{C}\wedge\omega^{C}{}_{B}$. (157)
Their spacetime components are related to the usual tensors by
$T^{A}{}_{\mu\nu}=e^{A}{}_{\rho}T^{\rho}{}_{\mu\nu},\qquad R^{A}{}_{B\mu\nu}=e^{A}{}_{\rho}E_{B}{}^{\sigma}R^{\rho}{}_{\sigma\mu\nu}$. (158)
F.4 Teleparallel geometry
Teleparallel gravity [10, 12] is characterized by
$R^{A}{}_{B}=0,\qquad Q_{\mu\nu\rho}=0,\qquad T^{\mu}{}_{\nu\rho}\neq 0.$ (159)
The vanishing curvature implies that the spin connection is a pure gauge,
$\omega^{A}{}_{B\mu}=(\Lambda^{-1})^{A}{}_{C}\,\partial_{\mu}\Lambda^{C}{}_{B}$. (160)
In the Weitzenböck gauge, one may choose
$\omega^{A}{}_{B\mu}=0,$ (161)
so that all gravitational information is encoded in the torsion tensor. This formulation is dynamically equivalent to general relativity and underlies the teleparallel limit discussed in Appendix E.
F.5 Bianchi identities
In differential-form notation, the Bianchi identities take the compact form
$DT^{A}=R^{A}{}_{B}\wedge e^{B},\qquad DR^{A}{}_{B}=0.$ (162)
In the teleparallel case, where $R^{A}{}_{B}=0$, this reduces to
$DT^{A}=0.$ (163)
These relations are the exterior calculus counterparts of the component Bianchi identities derived in Appendix D.
This appendix clarifies how the same spacetime geometry may be described equivalently in terms of curvature (Riemannian geometry), torsion (teleparallel geometry), or mixed curvature–torsion structures (metric–affine geometry). The equivalence underlying the wave equations derived in the main text supports the central theme of this work: geometry itself propagates.
APPENDIX G: HISTORICAL AND CONCEPTUAL DEVELOPMENT OF THE CURVATURE, THE TORSION, AND THE TELEPARALLEL GEOMETRIES
This appendix provides a historical and conceptual overview of the geometric ideas underlying the main body of the paper. Our aim is not merely historical completeness, but to clarify why the sequence of curvature $\rightarrow$ Ricci identities $\rightarrow$ torsion and nonmetricity $\rightarrow$ tetrads and spin connections $\rightarrow$ teleparallel formulations is both natural and unavoidable.
H.1 Riemann, curvature, and the non-commutativity of derivatives
The modern concept of curvature originates in the seminal work of Bernhard Riemann (1854) [6], who introduced the idea that geometry is encoded in how distances change infinitesimally. Riemann already understood that curvature measures the obstruction to Euclidean behavior, but it was only later that this idea was given its algebraic formulation.
In contemporary language, curvature is most cleanly characterized by the non-commutativity of covariant derivatives:
$\nabla_{\mu},\nabla_{\nu}]V^{\rho}=R^{\rho}{}_{\sigma\mu\nu}V^{\sigma}.$ (164)
This viewpoint reveals curvature not merely as a tensor, but as a statement about the failure of parallel transport to be path-independent. The Jacobi identity for differential operators then guarantees the existence of nontrivial identities satisfied by the curvature tensor, foreshadowing the Bianchi identities.
H.2 Ricci, Levi-Civita, and the classical tensor calculus
The systematic development of tensor calculus was carried out by Gregorio Ricci-Curbastro and Tullio Levi-Civita [7] at the turn of the twentieth century. Their construction culminated in the introduction of the Levi-Civita connection, uniquely characterized by
$\nabla_{\lambda}g_{\mu\nu}=0,\qquad T^{\rho}{}_{\mu\nu}=0.$ (165)
Within this framework, the Riemann tensor acquires additional symmetries, and the Ricci identity and Bianchi identities take their familiar forms. These identities are not optional mathematical decorations: they arise directly from the algebraic structure of covariant differentiation and ultimately encode the geometry’s internal consistency.
H.3 Einstein and the dynamical role of the Bianchi identities
In 1915, Albert Einstein elevated Riemannian geometry from a mathematical structure to a physical theory of gravitation [5]. By choosing the Levi-Civita connection [7], Einstein ensured that the contracted second Bianchi identity implies
$\nabla_{\mu}G^{\mu\nu}=0,$ (166)
which, in turn, enforces the covariant conservation of the energy-momentum tensor. Historically, this was a decisive insight: the Bianchi identities were recognized as the geometric reason why Einstein’s field equations are dynamically consistent. Thus, the curvature identities became the physical conservation laws.
H.4 Cartan geometry: torsion as an independent geometric field
Shortly after the formulation of general relativity, Elie Cartan [8] introduced a far-reaching generalization of geometry. In Cartan’s framework [8], the affine connection is no longer required to be symmetric or metric-compatible. Two new geometric objects emerge:
$T^{\rho}{}_{\mu\nu}=2\Gamma^{\rho}{}_{[\mu\nu]},\qquad Q_{\lambda\mu\nu}=\nabla_{\lambda}g_{\mu\nu}.$ (167)
Cartan reformulated geometry in terms of differential forms, introducing the coframe (tetrad) $e^{A}$ and the spin connection $\omega^{A}{}_{B}$. Curvature and torsion appear naturally as field strengths,
$T^{A}=De^{A},\qquad R^{A}{}_{B}=d\omega^{A}{}_{B}+\omega^{A}{}_{C}\wedge\omega^{C}{}_{B}$, (168)
and satisfy the generalized Bianchi identities. This formalism unifies the curvature, the torsion, and the nonmetricity into a single geometric language.
H.5 Spinors, tetrads, and the necessity of the spin connection
The introduction of spinor fields by Dirac in 1928 [21] revealed a crucial limitation of purely metric formulations of gravity. Spinors transform under local Lorentz transformations rather than general coordinate transformations, necessitating the introduction of tetrads and spin connections even within general relativity.
Thus, the tetrad–spin connection formalism is not an alternative description of gravity, but the natural framework once fermionic matter is included. Appendix F establishes this structure and shows how the spacetime tensors and the Lorentz-indexed objects are consistently related.
H.6 Teleparallel gravity and Einstein’s alternative viewpoint
In an attempt to unify gravity with electromagnetism, Einstein explored teleparallel or distant parallelism geometries in the late 1920s [10]. Although the unification program was abandoned, a key insight survived: the gravitational interaction can be encoded entirely in torsion, with vanishing curvature.
The modern teleparallel equivalent of general relativity (TEGR) is characterized by
$R^{\rho}{}_{\sigma\mu\nu}=0,\qquad Q_{\lambda\mu\nu}=0,\qquad T^{\rho}{}_{\mu\nu}\neq 0.$ (169)
Appendix E explicitly demonstrates how the Einstein–Hilbert scalar curvature can be rewritten as a torsion-squared invariant plus a boundary term. Thus, curvature and torsion provide equivalent descriptions of the same gravitational dynamics.
H.7 Symmetric teleparallel gravity and nonmetricity
A complementary formulation, known as symmetric teleparallel gravity, sets both curvature and torsion to zero while retaining nonmetricity:
$R^{\rho}{}_{\sigma\mu\nu}=0,\qquad T^{\rho}{}_{\mu\nu}=0,\qquad Q_{\lambda\mu\nu}\neq 0.$ (170)
In this case, gravity is entirely encoded in the failure of the connection to preserve the metric. This formulation completes a geometric triad:
Curvature Torsion Nonmetricity
$\neq 0$ =0=0 =0=0
=0=0 $\neq 0$ =0=0
=0=0 =0=0 $\neq 0$
The historical development reviewed here reveals a unifying theme:
  • Gravity can be understood as geometry, but the geometry itself admits multiple equivalent encodings.
  • Curvature, torsion, and nonmetricity represent different ways in which covariant derivatives fail to commute or preserve structure. The appendices of this paper make explicit how these descriptions are related and how general relativity emerges as a special, but not unique, geometric choice. For recent studies in various geometries, see [22, 23, 24, 25, 26, 27, 28] and the references therein.
Acknowledgements.
We thank Muzaffer Adak for useful discussions and for a careful reading of the manuscript. We are also grateful to Friedrich W. Hehl for an important remark that led to a revision of the table. Professor Metin Gürses has been a constant source of inspiration through his never-fading enthusiasm for physics, his exceptional work ethic, and his generosity as a collaborator and colleague. Having had the privilege of working with him for more than a decade on research and many joint papers, we deeply appreciate both his scientific insight and his kindness as a person. On the occasion of his becoming an emeritus professor at the Bilkent University Department of Mathematics, we celebrate his remarkable contributions to gravitational and mathematical physics and wish him many more years of joyful and fruitful exploration of gravity. E. Altas is supported by the TUBITAK Grant No. 123F353.